### v1

# Transformer models classify random numbers

Abstract
Random numbers are incredibly important in a variety of fields, and the need for their validation remains important. A Quantum Random Number Generator (QRNG) can theoretically generate truly random numbers however this does not remove the need to thoroughly test their randomness. Generally, the task of validating random numbers has been delegated to different statistical tests such as the tests from the NIST Statistical Test Suite (STS) which are often slow and only perform one task at a time. Our work presents a deep learning model that utilizes the transformer architecture to encode some of the tests from the NIST STS in a single model that also runs much faster. This model performs multi-label classification on these tests and outputs the probability of passing each statistical test that it encodes. We perform a thorough hyper-parameter optimization to converge on the best possible model and as a result, achieve a high degree of accuracy with a sample f1 score of above 0.9.
Keywords: Transformers, Random numbers, Multi-label classification
1 Introduction
Random numbers serve an important purpose in many fields, having numerous applications amongst each. Within cryptography, they are extensively studied and utilized to ensure secure encryption schemes keep our data safe [6, 25, 10, 7]. In physics, random numbers are highly studied in their appearances in quantum mechanics and thermodynamics [21, 3]. With random numbers having such a vast array of applications, obtaining and validating them also becomes a great issue[15]. Most of our encryption schemes today use pseudo-random numbers generated through pseudo-random number generators, and their validation has been delegated to a variety of statistical tests [1].
Along with using statistical tests for validating random numbers, deep learning and it’s derivatives can be used for determining the randomness of random numbers [8]. The uses of deep learning in this area have branched to to all of its facets, from Convolutional Neural Networks (CNNs), to Recurrent Neural Networks (RNNs), and Long Short Term Memory (LSTMs) which are types of RNNs [20, 26, 11].
More recently, the introduction of the transformer model has allowed further exploration of this deep learning method in the context of validating random numbers[27]. In the applications of the transformer models, the self-attention head is designed to detect sequential patterns[28]. Thus, in many applications where sequential data is available and easy to represent, the transformer has been quite successful, especially in the field of Natural Language Processing (NLP) but also in many others [13, 17, 22, 9, 4].
Since a random number by definition would have minimal sequential patterns and a non-random number would be the opposite, the self-attention mechanism in the transformer architecture should be quite effective at quantifying the randomness of random numbers. In the context of random numbers, traditionally, LSTMs were a popular deep learning method for qualifying the randomness of random numbers, but due to their complex nature as well as their inability to be parallelized, they were slow and often still inaccurate. As Li et. al. describe, these problems with LSTMs were solved by the transformer architecture [16]. Although the transformer architecture is well equipped for prediction of the next token, we focus on the quantification on the quality of randomness rather than what the next random number will be.
In this paper, we present a encoder-only transformer model that can accurately encode the statistical tests that were used quantify the quality of random numbers. The goal is to be able to encode multiple statistical tests (discussed further in the Methods section) into a single model that can be run once on a single binary sequence. We experiment with different architectures of this model as well as different hyper parameters to converge on the most optimized architecture for this problem. We show that by utilizing the transformer architecture, our model can accurately describe the type and degree of randomness of a binary sequence while improving efficiency over the statistical tests.
2 Methods
Non-deep learning based methods of validating random numbers include using a variety of statistical test suites such as the DIEHARD and DIEHARDER statistical test suites along with the NIST Statistical Test Suite (STS) [5, 18, 24]. These suites contain a multitude of tests that each test for a particular type of randomness. The purpose of this work is to encode a portion of these tests into a single transformer in order to test random numbers on multiple different tests at the same time.
2.1 Dataset
In general, we prepare the data to train our model by by first generating truly random numbers using a Quantum Random Number Generator (QRNG), augmenting it to have non random numbers as well as random numbers in the dataset, and then running the NIST STS to generate labels for our data. The purpose of the model is to run multi-label classification on given data so that the user can find out which tests the random number passes and which ones it fails. Thus, our model outputs a vector where each entry would be the probability of the random number passing the corresponding test. There were seven tests from the NIST STS that we encoded: Frequency, Block Frequency, Runs, Longest Run of Ones, Discrete Fourier Transform, Nonperiodic Template Matchings, Cumulative Sums. Out of the seven, we augmented data to generate non-random results for five of the tests: Frequency, Block Frequency, Runs, Longest Run of Ones, Nonperiodic Template Matching. In our findings we found that augmenting the data specifically tailored to these tests also resulted in the numbers failing the Discrete Fourier Transform and Cumulative Sums tests, which we did not augment the data for. Therefore, it was not necessary to augment the data to introduce non-randomness for these tests in order for the model to encode them.
2.2 Data Augmentation Techniques
The specific details of how these tests work are outline clearly by Ruhkin et al. in their definition of NIST STS[24]. The data augmentation techniques were adopted from Nagy et al. with negligible adjustments[20]
2.2.1 Frequency
We augment the frequency of either 0s or 1s in a stream by manually flipping a certain percentage of the bits to either 0 or one (depending on whether we are skewing the sequence to 0 or 1). A random number would have a roughly equal number of 0s and 1s. The aim of this test is to disrupt that balance.
2.2.2 Block Frequency
To augment a sequence to fail a block frequency test, we begin by dividng the sequence into M bit long chunks such that there are an even amount of evenly sized chunks, and then performing the frequency augmentation from above to each of these chunks. This preserves the overall ratio of 0s and 1s in the sequence, but within each chunk, the ratio is skewed to either 0 or 1. A random sequence broken into chunks would roughly preserve the ratio of half 0s and half 1s and the aim of this test is to disrupt that.
2.2.3 Runs
A ”run” is a sequence of like bits. A random sequence is full of small runs with larger runs becoming exponentially rarer the bigger they get. We augment the data to fail the runs test by offsetting the ratio of 0s and 1s in the sequence in chunks in an attempt to create fewer and larger runs, but keeping the overall ratio of 0s and 1s in the sequence constant. We do this by splitting the queue into M chunks with the same criteria as the Block Frequency test, and then offsetting the ratio of 0s and 1s in the chunk in the same way as the augmentation technique for the frequency test. We then scramble the runs to hide the subsequent checkerboard pattern.
2.2.4 Longest Run of Ones
To fail this test, instead of modifying the number of runs in the sequence, we aim to increase the length of the longest run. As mentioned before, it becomes exponentially rarer to get longer runs and so the longest run of a random sequence likely will not be too long. We again break the sequence into chunks as before, but instead of modifying the ratio of bits in each chunk, we insert a run of 1s and a run of 0s in the other to create long runs.
2.2.5 Nonperiodic Template Matching
The NIST STS provides a set of nonperiodic templates. The aim of this augmentation technique is to mimic a generator that produces the same binary sequences but not necessarily in a periodic fashion. A random number will feature the same binary sequences from time to time but they are quite rare. To train our model, we inserted 9 bit templates from the template set in each queue that was augmented to fail this test.
2.3 Model Training and Validation
Figure 1: Model Architecture. We trained 3 main models, one with 1 encoder layer, another with 3, and another with 4.
We begin by taking the binary sequence and tokenizing it to reduce its dimensionality. We trained many models to experiment with hyper-parameters which included tweaking the input size. We used inputs of size 1024 as well as 2048 bits to train our models. Our tokenization technique was simply getting the integer representation of every 16 bits. This would create a vocabulary for the transformer of size 65536 which was sufficiently large for model training. This means that our input would reduce to 64 or 128 tokens, which we would then run through the model whose architecture is shown above 1.
To monitor how well the model was training, we used aggregate f1 to measure model performance per batch as well as loss that best pertains to multi-label classification[2, 29]. We use a combination of sigmoid, binary cross entropy loss with the Adam optimizer with default parameters to train the model, and macro, micro, weighted, and sample f1 scores to validate it[14].
2.4 Handling Varying Input Size
One of the great perks of transformer models is their ability to handle data with no fixed input size [30]. However, as seen in Figure 1, since the output of the encoder is then flattened and connected into a fully connected layer, the input size must be fixed as a fully connected layer of an ANN cannot handle varying input size.
Figure 2: The effects on the shape of the input are illustrated. The averaging layer averages along each column to reduce the sequence length dimension to 1. The result is a vector that is the length of the embedding size. This vector is then passed into the fully connected layer as in Figure 1.
Our solution to this was quite simple. Since the fully connected layer can only take a fixed input length, we collapsed the only dimension that was varying which is the sequence length dimension. The drawback of this of course is that there is some information lost through the averaging process and how this affects model performance is discussed later.
3 Results
We set out with the goal of creating the best possible model for this task that can be used as an alternative to the statistical test suites. To find the model with the best hyper-parameters, we ran tests with different hyper-parameter settings to converge on the best possible model that we can find with our current architecture. The hyper-parameters that we tested for were the number of encoder layers, the size of the input, and the number of attention heads in the multi-headed attention section of the encoder. For brevity’s sake, we are only showing the sample f1 scores and omitting the other aggregate measures.
3.1 Encoder Layers
(a) Loss with 1 encoder layer
(b) Loss with 3 encoder layers
(c) Loss with 6 encoder layers
(d) F1 with 1 encoder layer
(e) F1 with 3 encoder layers
(f) F1 with 6 encoder layers
Figure 3: Training curves for different models with varying numbers of encoder layers where loss and f1 scores are shown. There were three different encoder configurations tested: 1 layer, 3 layers and 6 layers. Each of these models have 8 attention heads and took a fixed input of 128 tokens.
As Huang et al. show, more transformer encoder layers should improve the model accuracy and in our case, f1 score [12]. While we did observe this to be true, there are some important caveats. Namely, the effects were only beneficial until a certain point; in our case, once the number of encoder layers exceeded 3, the model stopped learning and failed to converge as seen in Figure 3. A possible conclusion that can be drawn from this is that the model became too large for the input and ended up adding too much noise for any useful information to be perceptible in later layers.
3.2 Input Size
(a) Loss with a fixed token size of 64
(b) Loss with a fixed token size of 128
(c) Loss with a varying token size of 128
(d) F1 with a fixed token size of 64
(e) F1 with a fixed token size of 128
(f) F1 with a varying token size of 128
Figure 4: Training curves for different models varying input sizes where loss and f1 scores are shown. We tested inputs which were 64 tokens, 128 tokens, and varying of length (they included the averaging layer). Each of these models had 3 encoder layers and 8 self-attention heads.
Token size for transformer models has been studied extensively and we wanted to see how previous findings would compare to ours[23]. In this experiment, we had two main input sizes that were varied. We inputted binary strings of size 1024 bits and 2048 bits which equated to 64 and 128 token inputs respectively. 3 encoder layers was a stable amount of encoder layers that did not lead to problems when testing on multiple input lengths and so we chose that to be fixed. As seen in Figure 4 the size of the input did not really matter all that much as the model seemed to converge quite well and in fact even quicker when the input size was larger. However, the problems arose when we had to average the output from the transformer encoder as described in Section 2.4. Due to the information lost from averaging the output from the encoder, the model converged at quite a high loss and over-trained quite quickly. The f1 score curve was also quite choppy, indicating that the model was struggling to converge at all.
3.3 Number of Attention Heads
(a) Loss with 1 self-attention head
(b) Loss with 8 self-attention heads
(c) Loss with 24 self-attention heads
(d) F1 with 1 self-attention head
(e) F1 with 8 self-attention heads
(f) F1 with 24 self-attention heads
Figure 5: Training curves for different models with a varying number of self-attention heads where loss and f1 scores are shown. Each model had 3 encoder layers and was inputted a fixed 128 tokens
The number of attention heads in the multi-headed attention layer is important to finding the most optimized model as the number of attention heads can quite readily affect the performance of a model [19]. Therefore, we also experimented with the number of multi-headed attention layers to find the most optimized model. As seen in Figure 5, the number of heads did not play a large role in the performance of the model with 8 heads being barely better than 1 head if at all. However, there was a downfall to adding too many attention heads as the model completely failed to converge with 24 attention heads suggesting that there is a limit to how many attention heads can be added to the model. It is probably a good practice to not use more than 8 heads anyways since adding more parameters adds overhead and also seems unnecessary seeing as the improvement was marginal from 1 to 8 heads.
Appendix A Appendix
A.1 F1 Scores
Here are the equations used to define the f1 metrics which were used to validate our model.
Micro F1 Score:
[…]
Macro F1 Score:
[…]
Weighted F1 Score:
[…]
Sample F1 Score:
[…]